OpenStudy (nuccioreggie):

Which expression is equivalent to 2(m − 4) + 1? 2m − 7 2m − 3 2m + 9 2m + 10

OpenStudy (bloosweater):

C

OpenStudy (bloosweater):

2(m - 4) + 1 2m - 8 + 1 2 m - 7 Oops it was A, sorry haha

OpenStudy (nuccioreggie):

@iGreen

OpenStudy (nuccioreggie):

@dmndlife24

OpenStudy (nuccioreggie):

@Thehulk49

OpenStudy (anonymous):

Hold on

OpenStudy (nuccioreggie):

Ok

OpenStudy (anonymous):

First Let's simplify step-by-step. 2(m−4)+1 Distribute: =(2)(m)+(2)(−4)+1 =2m+−8+1 Combine Like Terms: =2m+−8+1 =(2m)+(−8+1) =2m+−7 Answer: =2m−7

OpenStudy (nuccioreggie):

Ok

OpenStudy (anonymous):

Did that help?